A 0.223 kg block of ice at 0 deg C is placed into 1.68 L of water at 21.6 deg C.

What is the final temperature after all of the ice has melted and the system reaches equilibrium, in deg C?
Answer #1

here,

the mass of ice , m1 = 0.223 kg

volume of water , V = 1.68 L

the mass of water , m2 = 1.68 kg

let the final equilibrium temperature be Tf

using principle of calorimetry

heat gained by ice = heat lost by water

m1 * ( Lf + Cw * ( Tf - 0)) = m2 * Cw * (21.6 - Tf)

0.223 * ( 334000 + 4186 * (Tf)) = 1.68 * 4186 * ( 21.6 - Tf)

solving for Tf

Tf = 9.72 degree C

the final equilibrium temperature is 9.72 degree C